# LIGHTWEIGHT ERROR MITIGATION STRATEGIES for POST-TRAINING N:M ACTIVATION SPARSITY in LLMS

### v1

@@ Abstract
The demand for efficient large language model (LLM) inference has intensified the focus on sparsification techniques. While semi-structured (N:M) pruning is well-established for weights, its application to activation pruning remains underexplored despite its potential for dynamic, input-adaptive compression and reductions in I/O overhead. This work presents a comprehensive analysis of methods for post-training N:M activation pruning in LLMs. Across multiple LLMs, we demonstrate that pruning activations enables superior preservation of generative capabilities compared to weight pruning at equivalent sparsity levels. We evaluate lightweight, plug-and-play error mitigation techniques and pruning criteria, establishing strong hardware-friendly baselines that require minimal calibration. Furthermore, we explore sparsity patterns beyond NVIDIA’s standard 2:4, showing that the 16:32 pattern achieves performance nearly on par with unstructured sparsity. However, considering the trade-off between flexibility and hardware implementation complexity, we focus on the 8:16 pattern as a superior candidate. Our findings provide both effective practical methods for activation pruning and a motivation for future hardware to support more flexible sparsity patterns. Our code is available here.
@@ 1 Introduction
The expanding capabilities of Large Language Models (LLMs) have driven enormous demand for efficient AI inference. A common rule of thumb states that inference serving speed for a dense model with $N$ parameters scales as $\propto 1/\sqrt{N}$ (Erdil, 2025). Accelerating inference typically involves either reducing numerical precision via quantization to mitigate memory bottlenecks and enable faster arithmetic (Shvetsov et al., 2024) or sparsification to reduce the number of parameters (Maximov et al., 2025). Sparsity improves LLM efficiency in two key ways: (1) by reducing computation, and (2) by reducing I/O traffic for transferring parameters between memory and compute units – a major bottleneck during inference.
Weights vs. Activations. While sparse weights and activations offer identical theoretical FLOPs (floating point operations count), their practical implications differ. Weight sparsity enables static compression, efficiently reducing model storage and memory bandwidth. However, its static nature often causes irreversible model degradation. In contrast, activation sparsity is dynamic and input-adaptive, preserving model capacity, which we demonstrate in our work.
Sparsity in LLMs. Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023), especially due to the ReLU activation function in MLP blocks (Mirzadeh et al., 2023). Even if a model was trained with a non-ReLU activation function, we can still perform relufication Song et al. (2024b) to induce sparsity. This makes activation sparsification a natural choice for enhancing the model’s performance.
Accelerating LLMs with sparse activations. Indeed, recent results demonstrated that activation sparsification supported by specialized software can increase the model’s vector by matrix multiplication during the LLM decoding stage up to $2\timestimes$ (Song et al., 2024b; a; Liu et al., 2024; Lee et al., 2024). The idea behind such approaches is relatively straightforward - if a vector has many zeroes, we can disregard corresponding columns in a weight matrix $W$ on the fly. However, as batch size grows beyond oneone, it becomes challenging to gain benefits (Shrestha et al., 2025). Moreover, most previous methods mainly rely on sparsity induced only by activation functions (Mirzadeh et al., 2023) in MLP blocks.
@@
Focus and Motivation.
Applying N:M sparsity patterns to activations, while conceptually straightforward, raises several key questions: (Q1) Selection Strategy: As with weight sparsification, the criterion for selecting which activations to prune is crucial for maintaining performance (Zhelnin et al., 2024). (Q2) Error Mitigation: Fine-tuning is a standard method for recovering performance lost to post-training sparsification, but can be prohibitively expensive or risk compromising model safety alignment (Kharinaev et al., 2025). Therefore, we focus on lightweight, plug-and-play and hardware-friendly approaches that require only minimal calibration data (e.g., WikiText) or no data at all to enhance performance. (Q3) Beyond 2:4 Patterns: While 2:4 is the only hardware-supported semi-structured pattern, we explore a broader range {2:4,4:8,8:16,16:32} to motivate future hardware design. While we cannot directly compare acceleration gains between various patterns, our work solely focuses on model performance.
Our Contributions.
This paper makes the following key contributions:
  • Activation vs. Weight Sparsity Superiority: We show that, at similar sparsity levels, activation pruning consistently outperforms weight pruning across four diverse LLMs: Llama2-7B, Llama3-8.1B, Qwen2.5-7B, and Gemma3-4B.
@@
  • Lightweight Error Mitigation and Selection Criteria: We comprehensively evaluate four plug-and-play error mitigation techniques, three evaluated for the first time in the context of activation sparsity. We also evaluate three pruning selection criteria. These methods include statistical criteria such as median shift and variance correction and provide strong, hardware-friendly baselines for activation sparsification.
[…]
2 Hardware and Comparison of 2:4 and 8:16 Sparsity
One of the primary aims of our work is to motivate the design of hardware for more flexible sparsity patterns. Currently, only NVIDIA GPUs support structured sparsity, specifically the 2:4 pattern. This pattern achieves  1.5–1.7x inference acceleration and  1.5x energy reduction for 7B models, primarily from a 2x reduction in memory bandwidth (Fang et al., 2024; Lin et al., 2023). At the same time, any N:M pattern beyond 2:4, including 8:16, can provide the same 2x bandwidth reduction but with greater flexibility. In the general N:MN{:}M format, 2:4 sparsity permits $\binom{4}{2}=6$ non-zero patterns per block, requiring $\log_{2}(6)/4\approx 0.75$ bits per element for metadata. In contrast, the 8:16 pattern allows $\binom{16}{8}=12{,}870$ distinct configurations, demanding $\log_{2}(12{,}870)/16\approx 0.875$ bits per element. Stacking four 2:4 blocks to form a 16-element unit yields only $6^{4}=1{,}296$ configurations—nearly an order of magnitude fewer than the 12,870 offered by a native 8:16 block. Although 8:16 requires slightly more metadata storage, its vastly greater configurational flexibility makes it a compelling candidate. The main trade-off is implementation complexity. Decoding sparsity metadata and gathering non-zero values requires sophisticated circuitry. As the block size $M$ grows, these gather operations could impact performance. However, operating on larger, aligned blocks could also improve cache utilization, potentially amortizing this overhead.
@@ 2 Hardware and Comparison of 2:4 and 8:16 Sparsity
3 Methods

### v2

# Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches

@@ Abstract
[…]
Motivating Next-Gen Accelerators with Flexible $N:M$ Activation Sparsity via Benchmarking Lightweight Post-Training Sparsification Approaches
Shirin Alanova 1, Kristina Kazistova 1, Ekaterina Galaeva 1, Alina Kostromina 1, Vladimir Smirnov 1, Dmitry Redko 1, Alexey Dontsov 1, Maxim Zhelnin 1, Evgeny Burnaev 1,2, Egor Shvetsov 1 1 Skolkovo Institute of Science and Technology 2 Artificial Intelligence Research Institute 3 HSE University indicates equal contribution.
@@ 1 Introduction
Large Language Models (LLMs) have intensified demand for efficient inference. A common rule of thumb suggests serving speed for a dense $N$-parameter model scales as $\propto 1/\sqrt{N}$ (Erdil, 2025). Inference is often accelerated via quantization (Frantar et al., 2022; Shvetsov et al., 2024) or sparsification (Frantar and Alistarh, 2023; Maximov et al., 2025), with sparsity reducing both compute and memory I/O.
Weights vs. Activations. Although weight and activation sparsity yield the same theoretical FLOP count, they differ in practice. Weight sparsity enables static compression but can irreversibly degrade model quality, whereas activation sparsity is input-adaptive and better preserves model capacity though it requires an additional pruning step for each input.
Accelerating LLMs with Sparse Activations. Naturally, pruning is most effective when the values to be pruned already have low magnitudes or are zero, which is often the case for some LLMs’ intermediate representations (Liu et al., 2023; Li et al., 2022). Activation sparsity was employed to accelerate the decoding stage by enabling faster sparse vectordense matrix multiplications up to 2× speedup using specialized kernels (Song et al., 2024b, a; Liu et al., 2024; Lee et al., 2024). These gains are most pronounced for batch size 1 and diminish as batch size increases (Shrestha et al., 2025). The speedup primarily arises because rows of the dense matrix corresponding to zero elements in the sparse vector can be skipped during computation. Moreover, these methods often require some predictive mechanism to upload required weight indices into memory ahead of time Liu et al. (2023).
Semi-structured N:M activation sparsity, keeping $N$ non-zeros per block of size $M$, can extend benefits beyond single vector–matrix products by employing hardware support. Although 2:4 activation sparsity has been explored for training (Haziza et al., 2025; Wang et al., 2024), and post-training weight sparsification Maximov et al. (2025); Han et al. (2015); Frantar and Alistarh (2023); Kurtić et al. (2023) post-training N:M activation pruning is largely unexplored because it is not supported on commercial hardware.
@@
Focus and Motivation. The primary goal of this work is twofold: to motivate the development of hardware that natively supports semi-structured activation sparsity, and to benchmark existing approaches for inducing such sparsity in activations. Generally, these approaches address two key challenges:
(P1) Selection Strategy. As with weight sparsification, the choice of which activations to retain critically affects model accuracy (Zhelnin et al., 2025).
(P2) Error Mitigation. While post-training fine-tuning can recover lost performance, it is often impractical due to computational cost or risks to safety alignment (Kharinaev et al., 2025). We therefore mainly focus on lightweight, plug-and-play methods that require minimal (e.g., WikiText) or zero calibration data.
Our Contributions:
  • First, we demonstrate that activation sparsity outperforms weight sparsity. At matched sparsity levels, activation pruning consistently yields higher accuracy than weight pruning across four diverse large language models Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct highlighting activations as a more promising target for future sparse accelerators.
@@
  • Second, we benchmark four plug-and-play error mitigation techniques, three of which are applied to semi-structured activation sparsity for the first time, including statistical approaches such as median shift and variance correction. We also evaluate three selection criteria. These methods establish strong, retraining-free baselines that require minimal metadata, aligning well with hardware constraints.
  • Finally, we analyze a wide range of structured sparsity patterns and show that larger patterns dramatically improve model fidelity: the 16:32 pattern achieves performance close to unstructured 50% sparsity and retains over three times more accuracy than the conventional 2:4 pattern. Nevertheless, considering implementation trade-offs, we advocate for 8:16 as the optimal balance it offers twice the accuracy retention of 2:4 while remaining highly practical for near-term hardware adoption.
Together, these results provide concrete evidence that expanding hardware support beyond static 2:4 weight sparsity can unlock significant efficiency gains without compromising model quality, thereby motivating the next generation of sparsity-aware architectures.
2 Methods
In the sections below, we formally define the pruning criteria and error mitigation strategies evaluated in this work. A brief summary of these methods is provided in Table 1 and more detailed Table is available in Appendix 6.
Table 1: Summary of evaluated pruning metrics and transformations. Full descriptions are provided in Appendix 6.
Short name Description
Pruning metrics
ACT Activation magnitude
WT Weight magnitude
CLACT* Context-aware cosine-loss-inspired
Amber-Pruner Weight-informed importance
Transformations
D-PTS* Dynamic per-token shift
S-PTS Static per-token shift
L-PTS* Learnable per-token shift
VAR* Variance correction
VAR+L-PTS* VAR + learnable shift
R-Sparse Sparse activations + low-rank weights
2.1 Preliminaries
For a linear layer with weights $\mathbf{W}$, input activations $\mathbf{X}$, and output $\mathbf{Y}$:
$\mathbf{Y}=\mathbf{X}\mathbf{W}^{\top}.$ (1)
We construct a sparsity mask $\mathbf{M}$ using a metric $S$ and threshold $t$:
$\mathbf{M}_{ij}=\begin{cases}1,&S(\mathbf{X}_{ij})\geq t,\\ 0,&S(\mathbf{X}_{ij})<t,\end{cases}$ (2)
yielding
$\mathbf{Y}_{p}=(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}.$ (3)
Unstructured sparsity applies a global threshold over all elements. Semi-structured N:MN{:}M sparsity partitions rows (or columns) into non-overlapping blocks of size $M$ and keeps the top-$N$ elements per block by $S$ (e.g., 2:4 removes 50%) (Hu et al., 2024). We also study 8:16, which retains the same 50% density with higher flexibility and modest metadata overhead (Maximov et al., 2025).
2.2 Pruning Criterion
CLACT: Inspired by output cosine similarity in (Mi et al., 2025), we propose Cosine Loss ACTivation (CLACT), a context-aware score that emphasizes activations aligned with their row/column energy:
where $h$ is the hidden dimension and $l$ is the sequence length.
Amber-Pruner: Following (An et al., 2025), we (i) remove weight outliers outside the 0.5–99.5 percentiles, (ii) standardize the remaining weights, and (iii) score each activation as $S_{\textit{Amb.-Pr.}}(\mathbf{X}_{ij})=|\mathbf{X}_{ij}|\cdot\mathcal{L}(\hat{\mathbf{W}}_{:,j})$, where $\mathcal{L}(\cdot)$ is the channel-wise $\ell_{2}$ norm.
CLACT is context-aware (and for l=1l{=}1 reduces to an $\ell_{1}$-type criterion), whereas Amber-Pruner leverages weight magnitudes but is not context-aware.
2.3 Transformations for Error Mitigation
D-/S-/L-PTS (dynamic/ static/ learnable per-token shift) centers activations near zero: $\hat{\mathbf{X}}=\mathbf{X}-\bm{\eta}$, and uses the compensated form $\mathbf{Y}_{p}=((\hat{\mathbf{X}}\odot\mathbf{M})+\bm{\eta})\mathbf{W}^{\top}$ (Chua et al., 2024), where $\bm{\eta}=\overline{\mathbf{X}}$ for D-PTS; S-PTS uses a fixed $\bm{\eta}$ collected in a short warm-up; L-PTS learns $\bm{\eta}$.
VAR applies per-token variance correction after pruning: $\mathbf{Y}_{p}=\bm{\nu}(\mathbf{X}\odot\mathbf{M})\mathbf{W}^{\top}$, with
$\nu=\sqrt{\dfrac{\operatorname{Var}[\mathbf{X}]}{\operatorname{Var}[\mathbf{X}\odot\mathbf{M}]}}.$ (5)
VAR+L-PTS combines VAR scaling with the learnable shift. R-Sparse combines activation sparsity with a low-rank weight approximation (Kamirul et al., 2025) (details in Appendix B).
For methods requiring calibration/learning, we use WikiText-2: S-PTS stores a fixed bias vector $\bm{\eta}$, L-PTS (and VAR+L-PTS) learns $\bm{\eta}$, and R-Sparse learns low-rank factors; all other methods require no learned parameters.
2.4 Evaluation & Models
We evaluate the methods in Table 1 in two stages. First, we use Core Datasets (BoolQ, WinoGrande, PIQA, ARC-Easy) to screen all methods. We then focus on the most promising approaches on Qwen2.5-7B-Instruct and Llama3.1-8B-Instruct, and extend evaluation to Extended Datasets (HellaSwag, OpenBookQA, RTE, MMLU, Lambada_standard, Lambada_openai, IFEval). When calibration is required, we use WikiText-2, following common compression practice Egiazarian et al. (2024); Frantar et al. (2022); Van Baalen et al. (2024). All results are obtained with LM Eval Harness Gao et al. (2023); dataset details are in Table 9. We report results for Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct. For Qwen2.5-7B-Instruct, we do not sparsify key/query/value activations due to severe degradation observed in preliminary experiments.
@@ 2.4 Evaluation & Models
3 Results
3.1 Sparse Weights vs. Activations
3.2 Optimal semi-structured sparsity patterns
Our preliminary investigation demonstrates that while the 16:32 pattern achieves performance close to unstructured sparsity (a 5.4% drop versus 4.5% for 50% unstructured), it requires more metadata and greater resources for gather operations, as discussed in Section 1. Therefore, we focus on the 8:16 pattern, despite its higher performance drop of 7.38%. For comparison, the 2:4 pattern results in a 14.35% drop. These results are shown in Figure 2 and Table 7 in Appendix, we used only magnitude pruning to obtain these results. By demonstrating the superior model quality of 8:16 sparsity, our work incentivizes hardware designers to invest in or at least consider supporting the 8:16 pattern.
Table 2: Average relative performance drop (%, lower is better; negative = improvement) across four datasets, averaged over Llama2-7B-chat, Llama3.1-8B-Instruct, Qwen2.5-7B-Instruct and Gemma3-4B-Instruct. Act/Wt: activation/weight sparsity. Selection: ACT/CLACT/Amber-Pruner; transformations: VAR, D-PTS, S-PTS, L-PTS, R-SPARSE. Methods marked with an asterisk (*) are proposed in this paper. Full results: Appendix 11 and  12.
Target Pattern Method Avg drop ($\downarrow$)
Act u50 ACT $3.82$
Wt 2:4 WT $24.49$
Act 2:4 ACT $9.67$
Act 2:4 CLACT* $7.79$
Act 2:4 Amber-Pruner $7.85$
Act 2:4 VAR* $6.09$
Act 2:4 D-PTS* $5.84$
Act 2:4 S-PTS $4.29$
Act 2:4 L-PTS* $8.79$
Act 2:4 R-SPARSE (64) $7.70$
Act 2:4 R-SPARSE (128) $8.05$
Wt 8:16 WT $17.68$
Act 8:16 ACT $5.47$
Act 8:16 CLACT* $2.29$
Act 8:16 Amber-Pruner $1.56$
Act 8:16 VAR* $3.30$
Act 8:16 D-PTS* $2.07$
Act 8:16 S-PTS $0.61$
Act 8:16 L-PTS* $5.32$
Act 8:16 R-SPARSE (64) $1.52$
Act 8:16 R-SPARSE (128) $2.63$
3.3 Results on Single/Multi-choice Datasets
3.3.1 Evaluation of Pruning Criteria
We evaluate CLACT, Amber-Pruner, and magnitude pruning as a baseline. The main results for the 2:4 and 8:16 sparsity patterns are presented in Table 2. On average, both CLACT and Amber-Pruner outperform magnitude pruning by at least 2%, however, we observe no clear winner between them. As noted in Section 2.2, these criteria are designed for different purposes: CLACT adjusts based on context, while Amber-Pruner adjusts based on weight magnitudes. Notably, for Llama3.1-8B-Instruct under the 2:4 sparsity pattern, simple magnitude pruning outperforms both advanced criteria, underscoring model and architecture-specific sensitivities to pruning strategies.
3.3.2 Evaluation of Transformations
3.4 Instruction-Following Tasks
Table 3 presents instruction-following performance on the IFEval benchmark for Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, evaluated under two semi-structured sparsity patterns (2:4 and 8:16) and four activation transformation methods: S-PTS, D-PTS, R-Sparse, and VAR. First of all, we observe a strong model degradation on generative tasks. Second of all, we see that VAR is the strongest performer overall, especially for Llama3.1-8B-Instruct. S-PTS/D-PTS are competitive and lightweight, and R-Sparse lags significantly, particularly at 2:4. We speculate that while semi-structured patterns are good for prefill stage in LLMs they significantly degrade performance during decode stage. However, as discussed in Section 1 decode stage for single vector can be accelerated with more flexible approaches.
Table 3: Instruction-following (IFEval) prompt-level accuracy reported as PS/PL (prompt-level strict/loose accuracy).
Model Method 2:4 8:16
Llama3.1-8B ORIG 0.4455/0.4861 0.4455/0.4861
S-PTS 0.1682/0.1904 0.2995/0.3327
D-PTS 0.1941/0.2015 0.2828/0.3198
R-Sparse 0.0869/0.0979 0.2089/0.2311
VAR 0.2237/0.2458 0.3161/0.3586
Qwen2.5-7B ORIG 0.7135/0.7394 0.7135/0.7394
S-PTS 0.4325/0.5176 0.5194/0.5804
D-PTS 0.4399/0.5120 0.5434/0.5989
R-Sparse 0.2736/0.3457 0.3697/0.4196
VAR 0.4565/0.5342 0.5249/0.5896
3.5 Results for Unstructured Pruning
We evaluate D-PTS, VAR, and two selection criteria in this experiment: CLACT and Amber-Pruner using the Llama3.1-8B-Instruct model. Results in Table 4 indicate that VAR is the most effective transformation under unstructured sparsity. Moreover, CLACT outperforms Amber-Pruner by a wider margin here than in our semi-structured pruning experiments. These findings suggest two key insights: (1) No single method emerges as optimal for both unstructured and semi-structured sparsity. (2) The methods proposed in this work, VAR and CLACT, demonstrate strong generalization and are well-suited for both semi-structured and unstructured activation pruning.
Table 4: Performance comparison of pruning methods under 50% and 70% unstructured sparsity for Llama3.1-8B-Instruct model.
Method ArcE BoolQ PIQA Wino Grande (%) Avg. Drop
Original $0.821$ $0.839$ $0.800$ $0.734$
Unstructured 50%
ACT $0.777$ $0.820$ $0.771$ $0.686$ $4.450$
D-PTS $0.786$ $0.825$ $0.781$ $0.690$ $3.600$
VAR $0.784$ $0.819$ $0.776$ $0.705$ 3.470
CLACT $0.780$ $0.825$ $0.766$ $0.700$ $3.890$
Amber $0.768$ $0.820$ $0.763$ $0.702$ $4.450$
Unstructured 70%
ACT $0.558$ $0.631$ $0.647$ $0.548$ $25.320$
D-PTS $0.565$ $0.624$ $0.651$ $0.534$ $25.680$
VAR $0.614$ $0.651$ $0.676$ $0.532$ 22.660
CLACT $0.555$ $0.604$ $0.627$ $0.524$ $27.670$
Amber $0.487$ $0.594$ $0.590$ $0.539$ $30.680$
3.6 Combination of Methods
Table 5: Llama3.1-8B-Instruct with 8:16 activation sparsity: aggregated results across layer subsets. ORIG. AVG. = 0.6726. Drop is computed w/o perplexity. Full results: Appendix 13.
Method Layers PPL AVG. Drop $\downarrow$
LS+L-PTS all 9.6036 0.6047 10.90%
LS+L-PTS key,out,gate,down 8.3483 0.6385 5.43%
LS+L-PTS key,value,gate,down 8.0821 0.6503 3.56%
LS+L-PTS+VAR all 9.4983 0.6056 10.60%
LS+L-PTS+VAR key,out,gate,down 8.2930 0.6422 4.64%
LS+L-PTS+VAR key,value,gate,down 8.0259 0.6516 3.36%
3.7 Layer Sensitivity
We study layer sensitivity to 8:16 activation sparsity on Llama3.1-8B-Instruct using the Extended Datasets (Table 5). This analysis focuses on learnable methods, mainly L-PTS and LS (learnable diagonal scaling). Although learnable approaches underperform on average in our broader experiments, the best 8:16 configurations for Llama3.1-8B-Instruct rely on learnable parameters (Table 5). We observe that the FFN up projection and the attention out projection [^1] The output projection of the attention mechanism. It combines outputs from all attention heads and projects them back to the model’s hidden dimension. are the most sensitive: pruning them causes the largest drops, suggesting they should be preserved or handled with extra care. While this may not generalize to all layers, it indicates that layer importance under activation sparsity is highly non-uniform.
4 Discussion
5 Conclusion
This work establishes that post-training activation pruning is significantly more accuracy-preserving than weight pruning in large language models. Across four diverse architectures (Llama2-7B, Llama3.1-8B, Qwen2.5-7B, and Gemma3-4B), we demonstrate that activation sparsity consistently retains model capabilities better than weight sparsity at matched sparsity levels.
Our evaluation reveals that lightweight error mitigation techniques particularly CLACT, D-PTS, and VAR establish strong, hardware friendly baselines requiring minimal calibration data. Through systematic analysis of semi-structured patterns, we find that 16:32 approaches unstructured 50% sparsity in fidelity, while 8:16 emerges as the optimal near term target.
6 Limitations
Key limitations: First, all evaluations use software emulation without hardware measurements of speedup or energy efficiency. Second, layer sensitivity analysis remains preliminary; while FFN up-projections and attention out-projections appear most vulnerable, broader architectural studies are needed. Third, generative performance (IFEval) degrades significantly more than multiple-choice QA under sparsity, revealing an evaluation bias toward prefill-dominated workloads. The anomalous improvements on Qwen2.5-7B-Instruct for some benchmarks further highlight dataset-specific artifacts rather than genuine capability preservation. Our hardware overhead analysis in Appendix A is a rough estimate and more precise analysis is required.
Appendix
Appendix A Hardware Implications and Computational Overhead Analysis
While our empirical results demonstrate significant accuracy benefits from flexible N:M activation sparsity, the practical value of these techniques depends critically on whether hardware implementations can overcome the computational overhead of dynamic sparsification. This section provides a comprehensive break-even analysis to determine the hardware conditions required for activation sparsity to deliver net performance and efficiency gains.
The fundamental challenge for activation sparsity is that the theoretical memory bandwidth reduction must overcome the overhead introduced by dynamic sparsification operations. Based on literature and performance models, we estimate break-even thresholds where benefits begin to outweigh costs.
A.1 Energy-Delay Product Analysis
The Energy-Delay Product (EDP) provides a comprehensive metric for evaluating whether activation sparsity delivers net efficiency benefits. For 8:16 sparsity to be worthwhile from an EDP perspective, it must overcome both computational overhead and energy costs of the sparsification process itself.
We model EDP improvement for semi-structured sparsity patterns as:
$\text{EDP}_{\text{improvement}}=\frac{\text{EDP}_{\text{dense}}}{\text{EDP}_{\text{sparse}}}\approx\frac{r\cdot\eta}{1+\alpha}$
Where:
  • $r=2.0$ is the theoretical bandwidth reduction ratio for 8:16 sparsity
  • $\eta=0.85$ is the hardware utilization efficiency (representative of practical implementations)
  • $\alpha=0.3$ is the overhead factor from sparsification operations
This overhead factor $\alpha$ is calibrated from real measurements Fang et al. (2024) demonstrated that dynamic activation sparsification incurs 30–35% additional latency on current hardware without native support, broken down as:
  • Activation magnitude computation and block-wise selection
  • Mask application and metadata handling
  • Error mitigation techniques (D-PTS, VAR), which is not included in Fang et al. (2024).
Solving for the minimum hardware acceleration factor $k$ required for net EDP benefits:
$r\cdot\eta>k\cdot(1+\alpha)$
$2.0\cdot 0.85>k\cdot(1+0.3)$
$k>\frac{1.7}{1.3}\approx 1.31$
However, due to our imprecise estimations we will consider a higher amortized $k>1.6\times$ required for speedup.
A.2 Hardware Implementation Requirements
[…]
  • Dedicated sparsity controllers: On-chip circuitry that can generate sparsity masks with minimal latency, reducing the 15–20% selection overhead
  • Hardware-supported statistical units: Specialized units for variance/mean calculation required by error mitigation techniques, eliminating their computational overhead
  • Hierarchical sparsity support: Different patterns for different layer types based on sensitivity analysis (Section 3.7)
  • Bandwidth-optimized gather operations: Specialized memory controllers that maintain high efficiency despite irregular access patterns
Recent hardware developments show promising directions: NVIDIA’s Blackwell architecture includes a dedicated hardware decompression engine Jarmusch and Chandrasekaran (2025), while the LazyGPU microarchitecture enables lazy memory request issuing Liu et al. (2025).
[…]